OpenStudy (anonymous):

Find s10 for -1+-7+-13+-19+...

OpenStudy (solomonzelman):

what pattern do you notice in this sequence?

OpenStudy (anonymous):

the r=-6

OpenStudy (solomonzelman):

not, exactly. You go down by 6 every time (i.e. subtract 6) BUT the correct notation is d=-6

OpenStudy (solomonzelman):

when you say d=-6 that would mean you subtract -6, if you say r=-6 that would mean you are multiplying times -6 d = common difference (adding) r = common ratio (multiplying)

OpenStudy (solomonzelman):

You need to find the 10th term first, can you do that for me?

OpenStudy (solomonzelman):

[[ Use \(\large a_{\rm n}=a_1+{\rm d( n}-1) \) ]]

OpenStudy (solomonzelman):

r u lost?

OpenStudy (anonymous):

I got -55?

OpenStudy (solomonzelman):

Oh, a(10)=-1+(-6)(10-1)=-1+(-6)(9)=-1-54=-55 correct \(a_{10}=-55\)

OpenStudy (solomonzelman):

Now, (when you start from \(a_1\) and end the series at \(a_n\) (provided this series is arithmetic, which it is in this case) the sum, for n terms is given the following way: \(\large\color{black}{ \displaystyle \sum_{ {\rm n}=1 }^{ {\rm n} } A_{\rm n}=\color{red}{\frac{1}{2} \left(a_1+a_{\rm n}\right)}\times \color{blue}{{\rm n}} }\)

OpenStudy (solomonzelman):

in red I labeled the part of the formula which is the average term. in blue is the number of term

OpenStudy (solomonzelman):

number of terms*

OpenStudy (solomonzelman):

well, I should have made the number of terms k... but I will show you how to use this. \(\large\color{black}{ \displaystyle \sum_{ {\rm n}=1 }^{ {\rm 10} } A_{\rm n}=\color{red}{\frac{1}{2} \left(-1+-55\right)}\times \color{blue}{{\rm 10}} }\)

OpenStudy (solomonzelman):

in this case.... c y?

OpenStudy (solomonzelman):

(if you want I can restart in a more handy way, typing isn't a problem.... )

OpenStudy (anonymous):

I got -280 is that right? @SolomonZelman

OpenStudy (solomonzelman):

no